OpenStudy (anonymous):

Find the equation of a line with the given slope and y-intercept slope -5, y-intercept (0,-2)

sam (.sam.):

Put those into the point slope form \[y-y_1=m(x-x_1)\] Where m is -5 and \((x_1,y_1)\)

OpenStudy (anonymous):

I have solved it last night this was my finally answer Y=-2/5x-2 but I am not sure of it because am a bit confused ,can you check it for me

sam (.sam.):

Incorrect

sam (.sam.):

Try what I've told you to do, maybe you will get the answer

OpenStudy (anonymous):

Ok I will try and let you know what I got

sam (.sam.):

Alright

OpenStudy (anonymous):

I answered this... plug your slope in and find your y-intercept. y=mx+b

OpenStudy (anonymous):

That's not the way to solve it. The way you do it is putting it into the form y = mx+b, where b is the y-intercept and m is the slope. Thus, the equation is y = -5x - 2.

OpenStudy (anonymous):

What @whalexnuker said.

OpenStudy (anonymous):

y = -5x - 2 is correct.

OpenStudy (anonymous):

I am getting different final ,then whats is the final answer

OpenStudy (anonymous):

y = -5x - 2

OpenStudy (anonymous):

This was my final Y=-2/5x-2

OpenStudy (anonymous):

that's not correct. just use the equation y = mx +b, and plug the slope in for 'm', and the y-intercept in for 'b'.

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

Oh you are correct .I find out my mistake ,I got Y=-5x-2

OpenStudy (anonymous):

you could also do it the way sam said... y - y1 = m(x - x1) y - (-2) = -5(x - 0) y + 2 = -5x + 0 y = -5x - 2

OpenStudy (anonymous):

I plug in it the way same told me that is how I got my final answer ,Thanks guys

OpenStudy (anonymous):

no problem :)

OpenStudy (e.mccormick):

@Avihirschx Actually, .Sam.'s method is technically more correct. Not better, because it is slow, but more technically correct. If a person does not know that a y intercept of (0,2) means the same thing as b=2, then they have a conceptual issue. The proof to solve this conceptual issue is this: Given: Slope-Intercept form: \(y=mx+b\) Point-Slope form: \(y-y_1=m(x-x_1)\) The y intercept is where x is zero, \(\implies P=(0,y_1) \; \therefore x_1 = 0\) \(y-y_1=m(x-x_1)\implies \) \(y-y_1=m(x-0)\implies \) \(y-y_1=mx-0m\implies \) \(y-y_1=mx\implies \) \(y=mx+y_1 \) \(\therefore b= y_1\) when given the intercept in point form. Q.E.D. You already understand this and are able to quickly shift from point form to b. Great! However, the asker just did not get that. (I dealt with shakur for 2 hours yesterday on the same problem and was just looking to see if they ever got it straight.) Therefore, .Sam.'s method of having them put it in \(y-y_1=m(x-x_1)\) form might work and is mathematically using the steps that prove that \(b=y_1\) when it is given in the form of intercept \(= (0,y_1)\). That is what makes it more technically correct. Not better. I hope the person learns the relationship between the two. That way they can use the better, as in faster method.

OpenStudy (anonymous):

True. I intentionally submitted the easiest and quickest way to do that problem, not the most 'correct' way.

OpenStudy (e.mccormick):

Yah, which I tried yesterday... But it looks like shakur finally got it, so it is all good.